### v3

# Algebraic Hyperbolicity of Very General Hypersurfaces in Homogeneous Varieties

Abstract.
We generalize techniques by Coskun, Riedl, and Yeong, and obtain an almost optimal bound on the degree for the algebraic hyperbolicity of very general hypersurfaces in homogeneous varieties. As examples, we work out the cases of very general hypersurfaces in products of Grassmannians, orthogonal and symplectic Grassmannians, and flag varieties.
2020 Mathematics Subject Classification:
Primary: 32Q45, 14M17 ; Secondary: 14J70, 14M15
The author is supported by CNPq Post-doctoral fellowship 152039/2024-4
1. Introduction
A complex projective variety $X$ is algebraically hyperbolic if for some ample divisor $H$ there exists a real number $\epsilon>0$ such that for any integral curve $C\subset X$ the inequality
$2g(C)-2\geq\epsilon\deg_{H}(C)$
is satisfied, where $g(C)$ is the geometric genus of $C$. The goal of this paper is to prove the algebraic hyperbolicity of very general hypersurfaces in homogeneous varieties by generalizing techniques developed by Clemens [Cle86, Cle03], Ein [Ein88, Ein91], Voisin [Voi96, Voi98], Pacienza [Pac03, Pac04], Clemens and Ran [CR04], Coskun and Riedl [CR19, CR23] and Yeong [Yeo25].
A complex manifold $X$ is called (Brody) hyperbolic if every entire map $f:\mathbb{C}\to X$ is constant, and it is Kobayashi hyperbolic if the Kobayashi pseudometric is nondegenerate. In general, Kobayashi hyperbolicity implies Brody hyperbolicity, and when $X$ is a compact complex manifold, Brody proved that the converse holds [Bro78]. Hyperbolicity is conjectured to have deep connections to the geometry and arithmetic of varieties:
Conjecture 1.1.
(Lang’s Conjectures [Lan86])
  • A projective algebraic variety $X$ is hyperbolic if and only if every subvariety of $X$ is of general type.
  • A projective variety is hyperbolic if and only if it contains only finitely many rational points over any finite field extension.
Conjecture 1.2.
(Green-Griffiths-Lang Conjecture [GG80]) A smooth projective variety $X$ of general type contains a proper subvariety $S\subsetneq X$ containing all entire curves of $X$.
See Demailly’s survey paper [Dem20] for an exposition on hyperbolicity.
It can, however, be challenging to prove the hyperbolicity of a given variety. Demailly introduced algebraic hyperbolicity as an algebraic analogue for hyperbolicity, and proved that, for smooth projective varieties, hyperbolicity implies algebraic hyperbolicity and conjectured the converse [Dem97].
A lot of progress has been made in the study of the algebraic hyperbolicity of very general hypersurfaces in $\mathbb{P}^{n}$. For $n=3$, Xu [Xu94] proved that very general hypersurfaces $X\subset\mathbb{P}^{3}$ of degree $d\geq 6$ are algebraically hyperbolic. This was improved by Coskun and Riedl [CR19], who showed that a very general quintic surface is also algebraically hyperbolic. Since surfaces of degree at most four contain rational curves and cannot be algebraically hyperbolic, the classification for very general hypersurfaces in $\mathbb{P}^{3}$ is complete. For $n\geq 4$, Clemens and Ein [Cle86, Ein88] proved that $X$ is algebraically hyperbolic when $d\geq 2n$. This was improved to $d\geq 2n-1$ by Voisin [Voi96, Voi98]. Pacienza [Pac04] and Clemens and Ran [CR04] proved it for $d\geq 2n-2$ and $n\geq 6$. Yeong [Yeo25] did the octic 4-fold case and improved the result to $d\geq 2n-2$ and $n\geq 5$. If $d\leq 2n-3$, $X$ contains lines, thus it is not algebraically hyperbolic. Hence, the only remaining case in $\mathbb{P}^{n}$ is for sextic threefolds.
Haase and Ilten [HI21] started to classify algebraically hyperbolic surfaces in toric threefolds, and gave almost sharp bounds for the cases $\mathbb{P}^{2}\times\mathbb{P}^{1},\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$, the blowup of $\mathbb{P}^{3}$ at a point, and the weighted projective space $\mathbb{P}(1,1,1,n)$. Coskun and Riedl [CR23] streamlined their scroll method and used a careful analysis of where various line bundles fail to be section-dominating to obtain sharp bounds in the cases $\mathbb{P}^{2}\times\mathbb{P}^{1},\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$, the blowup of $\mathbb{P}^{3}$ at a point, and $\mathbb{F}_{e}\times\mathbb{P}^{1}$, and partial bounds for very general surfaces in other threefolds. Robins [Rob23] uses Haase and Ilten’s method to determine many algebraically hyperbolic surfaces in toric varieties with Picard rank 2 and 3. Yeong [Yeo25] applies Coskun and Riedl’s techniques to almost completely classify algebraically hyperbolic hypersurfaces in $\mathbb{P}^{m}\times\mathbb{P}^{n}$, except for some degrees in $\mathbb{P}^{3}\times\mathbb{P}^{1}$.
In this paper, we generalize the techniques of Coskun, Riedl, and Yeong to the context of more general homogeneous varieties. As an application, we obtain an almost sharp bound for the degree for which very general hypersurfaces of (products of) Grassmannians, orthogonal Grassmannians, symplectic Grassmannians, and flag varieties are algebraically hyperbolic.
Let us start by defining the general setting we will refer to throughout the paper.
Main Setting 1.3.
Let $A$ be a smooth complex projective variety such that:
  • $A$ is homogeneous with a transitive action by an algebraic group $G$;
  • $A$ has an embedding into a product of projective spaces
    $i:A\hookrightarrow\mathbb{P}^{N_{1}}\times\cdots\times\mathbb{P}^{N_{m}}$
    such that the divisors $H_{1},\ldots,H_{m}$ of $A$ corresponding to the pull-backs of the hyperplane classes $\mathcal{O}_{\mathbb{P}^{N_{i}}}(1)$ generate the Picard group of $A$;
  • The embedding $i$ is projectively normal, that is, the restriction
    $H^{0}\left(\mathbb{P}^{N_{1}}\times\cdots\times\mathbb{P}^{N_{m}},\mathcal{O}_{\mathbb{P}^{N_{1}}\times\cdots\times\mathbb{P}^{N_{m}}}(d_{1},\ldots,d_{m})\right)\to H^{0}(A,d_{1}H_{1}+\cdots+d_{m}H_{m})$
    is surjective for all $d_{1},\ldots,d_{m}>0$.
Write the class of the canonical divisor of $A$ as $K_{A}=a_{1}H_{1}+\cdots+a_{m}H_{m}$. Let $\mathcal{E}$ be the line bundle $d_{1}H_{1}+\cdots+d_{m}H_{m}$ for $d_{1},\ldots,d_{m}>0$, equivariant under $G$, and assume that $H_{1},\ldots,H_{m}$ is a collection of section-dominating line bundles for $\mathcal{E}$ (see Definition 2.3). A degree $(d_{1},\ldots,d_{m})$ hypersurface $X$ of $A$ is a section of $\mathcal{E}$.
Theorem 1.4.
Let $A$ be a homogeneous variety as in the Main Setting 1.3.
  • If $\dim A\geq 4$ and $d_{i}\geq\dim A-a_{i}-2$ for all $1\leq i\leq m$, then a very general hypersurface $X$ of degree $(d_{1},\ldots d_{m})$ is algebraically hyperbolic.
  • If $d_{i}\leq\dim A-a_{i}-4$ for some $1\leq i\leq m$, then a general hypersurface $X$ of degree $(d_{1},\ldots d_{m})$ contains lines. In particular, $X$ is not algebraically hyperbolic.
The cases with $d_{i}\geq\dim A-a_{i}-3$ for all $1\leq i\leq m$ and $d_{i}=\dim A-a_{i}-3$ for some $1\leq i\leq m$ remain open. When $\dim A\leq 4$, there are examples when it fails to imply algebraic hyperbolicity (see Examples 4.4 and 4.5.) For $A=\mathbb{P}^{4}$, the case of sextic threefolds remains open. However, when $\dim A\geq 5$, it does imply algebraic hyperbolicity for $A=\mathbb{P}^{n}$ and $A=\mathbb{P}^{m}\times\mathbb{P}^{n}$ [Yeo25]. We conjecture $X$ is algebraically hyperbolic in these cases when $\dim A\geq 5$ (see Conjecture 4.6.)
We then apply the theorem to concrete examples of homogeneous varieties. We note that the theorem can also be applied to products of these varieties.
Theorem 1.5.
Let $A$ be a:
  1. (1)
    Grassmannian $A=G(k,n)$
    • If $\dim A=k(n-k)\geq 4$ and $d\geq k(n-k)+n-2$, then a very general degree $d$ hypersurface of $G(k,n)$ is algebraically hyperbolic.
    • If $d\leq k(n-k)+n-4$, then a general degree $d$ hypersurface of $G(k,n)$ contains a line. In particular, it is not algebraically hyperbolic.
  2. (2)
    Product of Grassmannians $A=\prod_{i=1}^{m}G(k_{i},n_{i})$
    • If $\dim A=\sum_{i=1}^{m}k_{i}(n_{i}-k_{i})\geq 4$ and $d_{i}\geq\left(\sum_{j=1}^{m}k_{j}(n_{j}-k_{j})\right)+n_{i}-2$, for all $1\leq i\leq m$, then a very general hypersurface of $\prod_{i=1}^{m}G(k_{i},n_{i})$ of degree $(d_{1},\ldots,d_{m})$ is algebraically hyperbolic.
    • If $d_{i}\leq\left(\sum_{j=1}^{m}k_{j}(n_{j}-k_{j})\right)+n_{i}-4$ for some $i$, then a general hypersurface of degree $(d_{1},\ldots,d_{m})$ of $\prod_{i=1}^{m}G(k_{i},n_{i})$ contains a line. In particular, it is not algebraically hyperbolic.
  3. (3)
    Orthogonal Grassmannian $A=OG(k,n)$
    • If $\dim OG(k,n)=\frac{k(2n-3k-1)}{2}\geq 4$ and $d\geq\frac{k(2n-3k-1)}{2}+n-k-3$, then a very general hypersurface of $OG(k,n)$ of degree $d$ is algebraically hyperbolic.
    • If $d\leq\frac{k(2n-3k-1)}{2}+n-k-5$, then a general hypersurface of degree $d$ of $OG(k,n)$ contains a line. In particular, it is not algebraically hyperbolic.
  4. (4)
    Symplectic Grassmannian $A=SG(k,n)$
    • If $\dim A=\frac{k(2n-3k+1)}{2}\geq 4$ and $d\geq\frac{k(2n-3k+1)}{2}+n-k-1$, then a very general degree $d$ hypersurface of $SG(k,n)$ is algebraically hyperbolic.
    • If $d\leq\frac{k(2n-3k+1)}{2}+n-k-3$, then a general degree $d$ hypersurface of $SG(k,n)$ contains a line. In particular, it is not algebraically hyperbolic.
  5. (5)
    Flag Variety $A=F(k_{1},\ldots,k_{m};n)$
    • If $\dim A=\sum_{i=1}^{m}k_{i}(k_{i+1}-k_{i})\geq 4$ and $d_{i}\geq\left(\sum_{j=1}^{m}k_{j}(k_{j+1}-k_{j})\right)+k_{i+1}-k_{i-1}-2$ for all $1\leq i\leq m$, then a very general hypersurface of $F(k_{1},\ldots,k_{m};n)$ of degree $(d_{1},\ldots,d_{m})$ is algebraically hyperbolic.
    • If $d_{i}\leq\left(\sum_{j=1}^{m}k_{j}(k_{j+1}-k_{j})\right)+k_{i+1}-k_{i-1}-4$ for some $i$, then a general hypersurface of $F(k_{1},\ldots,k_{m};n)$ of degree $(d_{1},\ldots,d_{m})$ contains a line. In particular, it is not algebraically hyperbolic.
Organization of the paper
In section 2, we describe the general setup we work on and obtain a first, weaker bound on algebraic hyperbolicity. Section 3 is dedicated to working out the scroll argument to homogeneous varieties and in proving the algebraic hyperbolicity statement in Theorem 1.4. In Section 4, we determine when general hypersurfaces in homogeneous varieties contain lines, and conclude the non-algebraically hyperbolic cases from Theorem 1.4. In Section 5, we apply the bound obtained in the previous sections to specific examples of homogeneous varieties.
Acknowledgements
This manuscript is part of my Ph.D. thesis at the University of Illinois at Chicago. I would like to thank my advisor, Izzet Coskun, for our weekly meetings and support in writing this paper. I am grateful to Eric Riedl, Wern Yeong, and Jackson Morrow for their valuable advice and suggestions. I also want to thank the journal referees for identifying some mistakes and suggesting improvements.
2. Preliminaries
2.1. Setup
We first describe the general setup for the techniques developed by Clemens and Ein [Ein88, Ein91], Voisin [Voi96, Voi98], Pacienza [Pac03, Pac04], Coskun and Riedl [CR19, CR23], and Yeong [Yeo25].
Let $A$ be a smooth, complex projective variety of dimension $D$, and assume $A$ admits a transitive group action by an algebraic group $G$. Let $\mathcal{E}$ a globally generated vector bundle on $A$, equivariant under $G$, of rank $r<D-1$, and let $V=H^{0}(A,\mathcal{E})$. We will be particularly interested when $\mathcal{E}$ is a very ample line bundle on $A$.
Let $X$ be the zero locus of a very general section of $\mathcal{E}$, and suppose $X$ contains a curve $Y$ of degree $e$ and geometric genus $g$. Then, if $\mathcal{X}_{1}$ is the universal hypersurface over $V$, we get the relative Hilbert scheme $\mathcal{H}\to V$ with universal curve $\mathcal{Y}_{1}$, where the general fiber of $\mathcal{Y}_{1}\to\mathcal{H}$ is a curve of geometric genus $g$ and degree $e$. We can find a $G$-invariant subvariety $U\subset\mathcal{H}$ such that the map $U\to V$ is étale. Restricting $\mathcal{Y}_{1}$ to $U$ and taking a resolution of the general fiber, we get a smooth family $\mathcal{Y}\to U$ whose fibers are smooth curves of genus $g$. We pullback $\mathcal{X}_{1}$ to a family $\mathcal{X}$ over $U$, with projection maps $\pi_{1}:\mathcal{X}\to U$ and $\pi_{2}:\mathcal{X}\to A$. There is a natural generically injective map $h:\mathcal{Y}\to\mathcal{X}$.
We define the vertical tangent sheaf $T_{\mathcal{X}/A}$ by the short exact sequence
$0\longrightarrow T_{\mathcal{X}/A}\longrightarrow T_{\mathcal{X}}\longrightarrow\pi_{2}^{*}T_{A}\longrightarrow 0$
Since we constructed $\mathcal{Y}$ to be stable under the $G$-action, we have that $\pi_{2}\circ h$ dominates $A$ and that the map $T_{\mathcal{Y}}\to h^{*}\pi_{2}^{*}T_{A}$ is surjective. Similarly, we define the vertical tangent sheaf $T_{\mathcal{Y}/A}$ as the kernel in the short exact sequence
$0\longrightarrow T_{\mathcal{Y}/A}\longrightarrow T_{\mathcal{Y}}\longrightarrow h^{*}\pi_{2}^{*}T_{A}\longrightarrow 0$
Let $M_{\mathcal{E}}$ denote the Lazarsfeld-Mukai bundle associated to $\mathcal{E}$, defined by the sequence
$0\longrightarrow M_{\mathcal{E}}\longrightarrow H^{0}(A,\mathcal{E})\otimes\mathcal{O}_{A}\longrightarrow\mathcal{E}\longrightarrow 0$
For a general element $t$ of $U$, let $Y_{t}$ be the fiber of $\mathcal{Y}$ over $t$ and $X_{t}$ be the fiber of $\mathcal{X}$ over $t$. Denote by $h_{t}:Y_{t}\to X_{t}$ the restriction of $h$ to $Y_{t}$. The following result translates the relation between the vertical tangent sheaves and the Lazarsfeld-Mukai bundle.
Proposition 2.1.
(cf. [CR23, Proposition 2.1])
  1. (1)
    $N_{h_{t}/X_{t}}\cong N_{h/\mathcal{X}}|_{Y_{t}}$.
  2. (2)
    $T_{\mathcal{X}/A}\cong\pi_{2}^{*}M_{\mathcal{E}}$.
  3. (3)
    $N_{h/\mathcal{X}}$ is the cokernel of the map of vertical tangent sheaves $T_{\mathcal{Y}/A}\to h^{*}T_{\mathcal{X}/A}$.
2.2. The degree of the normal bundle and section-dominating line bundles
There is a close relation between the degree of the normal bundle $N_{h_{t}/X_{t}}$ and the genus of the curve $Y_{t}$ described by the following lemma. We can use it to bound the genus of $Y_{t}$ by finding a bound for the degree of the normal bundle $N_{h_{t}/X_{t}}$.
Lemma 2.2.
(cf. [CR23, Lemma 2.2]) The following holds:
$\deg(N_{h_{t}/X_{t}})=2g(Y_{t})-2-K_{X_{t}}\cdot h_{t}(Y_{t}).$
By Proposition 2.1, we have a surjection
(2.1) $M_{\mathcal{E}}|_{Y_{t}}\to N_{h_{t}/X_{t}}.$
We can get a surjection to $N_{h_{t}/X_{t}}$ from a simpler bundle by using the notion of section-dominating line bundles.
Definition 2.3.
(cf. [CR23, Definition 2.3]) Let $\mathcal{E}$ be a vector bundle on $A$. A collection of non-trivial globally generated line bundles $L_{1},\ldots,L_{m}$ is a section-dominating collection of line bundles for $\mathcal{E}$ if $\mathcal{E}\otimes L_{i}^{\vee}$ is globally generated for every $1\leq i\leq m$ and the map
$\bigoplus_{i=1}^{m}\left(H^{0}(L_{i}\otimes\mathcal{I}_{p})\otimes H^{0}(\mathcal{E}\otimes L_{i}^{\vee})\right)\to H^{0}(\mathcal{E}\otimes\mathcal{I}_{p})$
is surjective for every point $p\in A$.
Example 2.4.
Suppose $A=\mathbb{P}^{n}$ and let $H$ be the hyperplane class. Then, for any $\mathcal{E}=dH$ with $d>0$, $H$ is a section-dominating line bundle for $\mathcal{E}$. To prove this, we need to show that
$H^{0}(\mathcal{O}_{\mathbb{P}^{n}}(1)\otimes\mathcal{I}_{p})\otimes H^{0}(\mathcal{O}_{\mathbb{P}^{n}}(d-1))\to H^{0}(\mathcal{O}_{\mathbb{P}^{n}}(d)\otimes\mathcal{I}_{p})$
is surjective. Choose coordinates $x_{0},\ldots,x_{n}$ for $\mathbb{P}^{n}$ so that $p=(1:0:\cdots:0)$. Then $H^{0}(\mathcal{O}_{\mathbb{P}^{n}}(d)\otimes\mathcal{I}_{p})$ is the set of degree $d$ forms without $x_{0}^{d}$ term, and these can be written as a combination of products of linear forms without $x_{0}$ term in $H^{0}(\mathcal{O}_{\mathbb{P}^{n}}(1)\otimes\mathcal{I}_{p})$ and degree $d-1$ forms in $H^{0}(\mathcal{O}_{\mathbb{P}^{n}}(d-1))$.
Example 2.5.
Let $A=G(k,n)$ be the Grassmannians of $k$-planes in the affine $n$-space, and let $H$ correspond to the hyperplane class, that is, the pullback $p^{*}\mathcal{O}_{\mathbb{P}^{N}}(1)$ via the Plücker embedding $p:G(k,n)\to\mathbb{P}^{N}$. Then, for any $\mathcal{E}=dH$, $d>0$, $H$ is a section-dominating line bundle for $\mathcal{E}$. This follows directly from the previous example and the projective normality of the Grassmannian under the Plücker embedding (cf. [RR85].) By [RR85], a similar result is true for flag varieties and Schubert varieties.
Example 2.6.
This property can be extended to products: let $A=\prod_{i=1}^{m}A_{i}$ be a product of varieties $A_{i}$ such that for each $i$, $H_{i}$ is a section-dominating line bundle for $\mathcal{E}_{i}=d_{i}H_{i}$ for some $d_{i}>0$. Then $H_{1},\ldots,H_{m}$ is a section-dominating collection of line bundles for $\mathcal{E}=\sum_{i=1}^{m}d_{i}H_{i}$.
Given a section of $H^{0}(\mathcal{E}\otimes L_{i}^{\vee})$, the natural multiplication map defines maps $L_{i}\to\mathcal{E}$ and $H^{0}(L_{i})\to H^{0}(\mathcal{E})$, and induces a map $M_{L_{i}}\to M_{\mathcal{E}}$. By choosing bases for $H^{0}(\mathcal{E}\otimes L_{i}^{\vee})$, we can get a surjection to $M_{\mathcal{E}}$.
Proposition 2.7.
(cf. [CR23, Proposition 2.7]) Let $\mathcal{E}$ be a globally generated vector bundle and $M_{\mathcal{E}}$ the Lazarsfeld-Mukai bundle associated to $\mathcal{E}$. Let $L_{1},\ldots,L_{m}$ be a section-dominating collection of line bundles for $\mathcal{E}$. Then, there is a surjection
$\bigoplus_{i=1}^{m}M_{L_{i}}^{\oplus s_{i}}\to M_{\mathcal{E}}$
for some integers $s_{i}$.
From Proposition 2.7 and the surjection (2.1), we obtain a surjection to the normal bundle
(2.2) $\beta:\bigoplus_{i=1}^{m}M_{L_{i}}^{\oplus s_{i}}|_{Y_{t}}\to N_{h_{t}/X_{t}}.$
By choosing generic sections of $H^{0}(\mathcal{E}\otimes L_{i}^{\vee})$ for the multiplication in Proposition 2.7, the rank of the image of $\beta$ is a strictly increasing function of the $s_{i}$ until surjectivity is reached (cf. [Cle03, Section 2.2].) In particular, we can choose a surjection $\beta$ so that $s_{1}+\cdots+s_{m}\leq\operatorname{rk}N_{h_{t}/X_{t}}$.
This surjection can be used to get a bound for the degree of $N_{h_{t}/X_{t}}$ in terms of the integers $s_{i}$.
Lemma 2.8.
Let $\beta:\bigoplus_{i=1}^{m}M_{L_{i}}^{\oplus s_{i}}|_{Y_{t}}\to N_{h_{t}/X_{t}}$ be a surjective map as in (2.2). Then
$\deg N_{h_{t}/X_{t}}\geq-\sum_{i=1}^{m}s_{i}\deg L_{i}|_{Y_{t}}.$
Proof.
Let $K$ be the kernel of $\beta$:
[…]
Then by the definition of the $M_{L_{i}}$, we get an injection $K\to\bigoplus_{i=1}^{m}H^{0}(A,L_{i})^{\oplus s_{i}}\otimes\mathcal{O}_{A}|_{Y_{t}}$. Since it is a direct sum of trivial bundles, $\bigoplus_{i=1}^{m}H^{0}(A,L_{i})^{\oplus s_{i}}\otimes\mathcal{O}_{A}|_{Y_{t}}$ is a semi-stable vector bundle. It follows that
$\frac{\deg K}{\operatorname{rk}K}=\mu(K)\leq 0,$
thus
$\deg(K)\leq 0,$
hence, from the sequence above and the sequence defining $M_{L_{i}}$, we have
$\deg N_{h_{t}/X_{t}}=\deg\left(\bigoplus_{i=1}^{m}M_{L_{i}}^{\oplus s_{i}}|_{Y_{t}}\right)-\deg K\geq\deg\left(\bigoplus_{i=1}^{m}M_{L_{i}}^{\oplus s_{i}}|_{Y_{t}}\right)=-\sum_{i=1}^{m}s_{i}\deg L_{i}|_{Y_{t}}.$
By combining Lemmas 2.2 and 2.8, we obtain a bound for the genus:
(2.3) $2g(Y_{t})-2\geq K_{X_{t}}\cdot Y_{t}-\sum_{i=1}^{m}s_{i}\deg L_{i}|_{Y_{t}}.$
Therefore, if we can control the integers $s_{i}$, we can prove the algebraic hyperbolicity of $X_{t}$. This leads us to our first criterion for hyperbolicity. This result is similar to [CR23, Cor.2.9] and [HI21, Thm 3.6].
Proposition 2.9.
Let $A$ be a homogeneous variety of dimension $D$ and let $\mathcal{E}$ be a very ample line bundle on $A$ as in Setup 2.1. Let $L_{1},\ldots,L_{m}$ be a collection of section-dominating line bundles for $\mathcal{E}$ on $A$. If, for some $\epsilon>0$,
$K_{X_{t}}\cdot Y_{t}\geq(D-2+\epsilon)\cdot\sum_{i=1}^{m}\deg L_{i}|_{Y_{t}},$
then $X_{t}$ is algebraically hyperbolic.
Proof.
As before, we can choose the $s_{i}$ so that $s_{1}+\cdots+s_{m}\leq\operatorname{rk}N_{h_{t}/X_{t}}=D-2$. In particular, we can assume $s_{i}\leq D-2$ for every $1\leq i\leq m$. Thus, by inequality (2.3), we get
$2g(Y_{t})-2\geq K_{X_{t}}\cdot Y_{t}-(D-2)\sum_{i=1}^{m}\deg L_{i}|_{Y_{t}}.$
Therefore, if $K_{X_{t}}\cdot Y_{t}\geq(D-2+\epsilon)\cdot\sum_{i=1}^{m}\deg L_{i}|_{Y_{t}}$, we get
$2g(Y_{t})-2\geq\epsilon\cdot\sum_{i=1}^{m}\deg L_{i}|_{Y_{t}}.$
Hence, $X_{t}$ is algebraically hyperbolic. ∎
Remark 2.10.
This initial setup can also be defined when $A$ is not homogeneous but contains a Zariski-open homogeneous set $A_{0}$. Results equivalent to Proposition 2.9 can be obtained in this case, see for example, [CR23, Theorem 1.2]. This allowed Coskun and Riedl [CR23] to classify algebraically hyperbolic surfaces $X$ in the cases when $A$ is a Hirzebruch surface $\mathbb{P}(\mathcal{O}_{\mathbb{P}^{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}}(e))$, $A$ is $\mathbb{P}^{3}$ blown up at a single point, and $A$ is a weighted projective space $\mathbb{P}(1,1,1,n)$. We believe the generalizations made in this paper can also be worked out when $A$ contains a Zariski-open homogeneous set $A_{0}$, allowing future studies of higher-dimensional toric varieties, such as blowups of projective spaces and weighted projective spaces.
3. Scroll method for homogeneous varieties
In this section, we work out the scroll method introduced by Coskun and Riedl in [CR23] in a more general setting. This technique builds a surface scroll on the curve $Y_{t}$ whose degree can be used to improve the bound for algebraic hyperbolicity obtained in Proposition 2.9.
Here we will work on homogeneous varieties satisfying the Main Setting 1.3, which we recall here:
Let $A$ be a smooth complex projective variety of dimension $D$ such that:
  • $A$ is homogeneous with a transitive action by an algebraic group $G$,
  • $A$ has a projectively normal embedding into a product of projective spaces $\mathbb{P}^{N_{1}}\times\cdots\times\mathbb{P}^{N_{m}}$,
  • The divisors $H_{1},\ldots,H_{m}$ of $A$ corresponding to the pull-backs of the hyperplane classes $\mathcal{O}_{\mathbb{P}^{N_{i}}}(1)$ generate the Picard group of $A$.
Let $\mathcal{E}$ be the line bundle $d_{1}H_{1}+\cdots+d_{m}H_{m}$ for $d_{1},\ldots,d_{m}>0$, equivariant under $G$, and assume that $H_{1},\ldots,H_{m}$ is a collection of section-dominating line bundles for $\mathcal{E}$. To simplify the notation, we write $X$ for the zero locus of a very general section of $\mathcal{E}$, and $C$ for the curve in $X$, instead of $X_{t}$ and $Y_{t}$.
Write the class of the canonical divisor of $A$ as $K_{A}=a_{1}H_{1}+\cdots+a_{m}H_{m}$. By the adjunction formula, $K_{X}=\sum_{i=1}^{m}(a_{i}+d_{i})H_{i}|_{X}$. In this case, Proposition 2.9 can be rephrased as follows.
Corollary 3.1.
If there exists $\epsilon>0$ such that $d_{i}\geq(D-a_{i}-2)+\epsilon$ for all $1\leq i\leq m$, then $X$ is algebraically hyperbolic.
Proof.
Since
$K_{X}\cdot C=\sum_{i=1}^{m}(a_{i}+d_{i})H_{i}\cdot C\geq(D-2+\epsilon)\cdot\sum_{i=1}^{m}H_{i}\cdot C,$
the result follows from Proposition 2.9. ∎
The main theorem in this section shows that we do not need the $\epsilon$ in the bound above.
Definition 3.2.
Given a product of projective spaces $\mathbb{P}^{N_{1}}\times\cdots\times\mathbb{P}^{N_{m}}$, the fibers of the projection
$\pi_{1}^{\prime}:\mathbb{P}^{N_{1}}\times\cdots\times\mathbb{P}^{N_{m}}\to\mathbb{P}^{N_{2}}\times\cdots\times\mathbb{P}^{N_{m}}$
are $\mathbb{P}^{N_{1}}$. We define a $\mathbb{P}^{N_{1}}$-line as a line contained in one of these fibers. Equivalently, a $\mathbb{P}^{N_{1}}$-line is a curve of numerical class $H_{1}^{N_{1}-1}H_{2}^{N_{2}}\cdots H_{m}^{N_{m}}$. We define $\mathbb{P}^{N_{i}}$-lines similarly for $i=1,\ldots,m$. We call a surface scroll a $\mathbb{P}^{N_{i}}$-scroll if it contains a $\mathbb{P}^{N_{i}}$-line through every point.
We can see the space of $\mathbb{P}^{N_{1}}$-lines as the product $G(2,N_{1}+1)\times\mathbb{P}^{N_{2}}\times\cdots\times\mathbb{P}^{N_{m}}$ and the Fano scheme of $\mathbb{P}^{N_{1}}$-lines contained in $A$ as a subscheme $F_{1}^{N_{1}}(A)\subset G(2,N_{1}+1)\times\mathbb{P}^{N_{2}}\times\cdots\times\mathbb{P}^{N_{m}}$. Similarly for $\mathbb{P}^{N_{i}}$-lines. We will need the dimension of the space of $\mathbb{P}^{N_{i}}$-lines in $A$.
Lemma 3.3.
Let $F_{1}^{N_{i}}(A)$ be the Fano scheme of $\mathbb{P}^{N_{i}}$-lines in $A$. Then
$\dim F_{1}^{N_{i}}(A)=\dim A-a_{i}-3.$
Proof.
Let $L$ be a $\mathbb{P}^{N_{i}}$-line in $A$. Since $A$ is a homogeneous variety, its tangent bundle $T_{A}$ is globally generated, thus $T_{A}|_{L}$ is globally generated. A classic deformation argument (cf. [Deb01, Theorem 2.6]) then shows that $\dim F_{1}^{N_{i}}(A)=h^{0}(L,N_{L/A})=\dim A-K_{A}\cdot L-3=\dim A-a_{i}-3$. ∎
We use the following lemma to construct a scroll.
Lemma 3.4.
(cf. [CR23, Lemma 2.13]) A rank one quotient $Q$ of $M_{H_{i}}|_{C}$ induces a $\mathbb{P}^{N_{i}}$-scroll $\Sigma$ over $C$ of $H_{i}$-degree equal to $\deg Q+(H_{i}\cdot C)$.
Theorem 3.5.
Let $A$ be a homogeneous variety as in the Main Setting 1.3. If $\dim A\geq 4$ and $d_{i}\geq\dim A-a_{i}-2$ for all $1\leq i\leq m$, then $X$ is algebraically hyperbolic.
Proof.
By Proposition 2.7, there is a surjection
$\beta:\bigoplus_{i=1}^{m}M_{H_{i}}^{\oplus s_{i}}|_{C}\to N_{C/X}.$
Without loss of generality, assume $s_{1}\geq\cdots\geq s_{m}$. As before, we can assume $s_{1}+\cdots+s_{m}\leq\operatorname{rk}N_{C/X}=D-2$. We divide the proof into the cases $s_{1}\leq D-3$ and $s_{1}=D-2$.
Case 1. Suppose first that $s_{1}\leq D-3$. So, $s_{i}\leq D-3$ for all $i$. Then, the inequality
$2g(C)-2\geq K_{X}\cdot C-\sum_{i=1}^{m}s_{i}\deg H_{i}|_{C}$
from (2.3) implies that
$2g(C)-2\geq\left(\sum_{i=1}^{m}(a_{i}+d_{i})H_{i}\right)\cdot C-(D-3)\sum_{i=1}^{m}(H_{i}\cdot C)$
or equivalently,
$2g(C)-2\geq\sum_{i=1}^{m}(a_{i}+d_{i}-D+3)(H_{i}\cdot C).$
Since, by hypothesis, $d_{i}\geq D-a_{i}-2$ for $i=1,\ldots,m$, we get
$2g(C)-2\geq\sum_{i=1}^{m}(H_{i}\cdot C).$
Case 2. Suppose now instead that $s_{1}=D-2$ is the minimum possible value for $s_{1}$ for which $\beta$ is surjective. So we have $s_{2}=\cdots=s_{m}=0$ and a surjection
$\beta:M_{H_{1}}^{\oplus D-2}|_{C}\to N_{C/X}.$
Since this is the minimum value for $s_{1}$ such that $\beta$ is surjective, this means that the induced map
(3.1) $M_{H_{1}}|_{C}\to N_{C/X}/\beta\left(M_{H_{1}}^{\oplus D-3}|_{C}\right)$
has a rank-one image $Q$ (cf. [Cle03, Section 2.2].) By Lemma 3.4, this induces a $\mathbb{P}^{N_{1}}$-scroll $\Sigma\subset\mathbb{P}^{N_{1}}\times\cdots\times\mathbb{P}^{N_{m}}$ containing $C$ of $H_{1}$-degree $\Sigma\cdot H_{1}^{2}=\deg Q+(H_{1}\cdot C)$. And since $\Sigma$ is a $\mathbb{P}^{N_{1}}$-scroll, we have $H_{1}H_{j}\cdot\Sigma=H_{j}\cdot C$ for $j\neq 1$. Thus, $\Sigma$ has numerical class
$(H_{1}\cdot C+\deg Q)H_{1}^{N_{1}-2}H_{2}^{N_{2}}\cdots H_{m}^{N_{m}}+(H_{2}\cdot C)H_{1}^{N_{1}-1}H_{2}^{N_{2}-1}\cdots H_{m}^{N_{m}}+\cdots+(H_{m}\cdot C)H_{1}^{N_{1}-1}H_{2}^{N_{2}}\cdots H_{m}^{N_{m}-1}.$
Since $A$ is projectively normal, there exists a degree $(d_{1},\ldots,d_{m})$ hypersurface $Z\subset\mathbb{P}^{N_{1}}\times\cdots\times\mathbb{P}^{N_{m}}$ such that $X=A\cap Z$. We divide into the cases $\Sigma\subset Z$ and $\Sigma\not\subset Z$.
Suppose first that $\Sigma\subset Z$. We will prove in Theorem 4.2 that $X$ does not contain $\mathbb{P}^{N_{1}}$-lines. Thus, we must have $\Sigma\not\subset A$. As projective homogeneous varieties in projective space are cut out by quadrics (cf. [Ram87, Theorem 3.11 and the following Remark 3.12]), by the Segre embedding we can express $A$ as cut out by varieties of degree (2,,2)(2$\ldots$,2)( 2 , … , 2 ) in $\mathbb{P}^{N_{1}}\times\cdots\times\mathbb{P}^{N_{m}}$.
Then, there exists a hypersurface $F$ of class $2H_{1}+\cdots+2H_{m}$ such that $\Sigma\not\subset F$. Intersecting the classes of $\Sigma$ and $F$, we have
$([\Sigma]\cdot[F])\cdot H_{1}=2(H_{1}\cdot C+\deg Q)+2(H_{2}\cdot C)+\cdots+2(H_{m}\cdot C).$
Since $C\subset\Sigma\cap F$, it follows that $([\Sigma]\cdot[F])\cdot H_{1}\geq H_{1}\cdot C$, so
$2(H_{1}\cdot C+\deg Q)+2(H_{2}\cdot C)+\cdots+2(H_{m}\cdot C)\geq H_{1}\cdot C$
hence
(3.2) $\deg Q\geq-\frac{1}{2}(H_{1}\cdot C)-(H_{2}\cdot C)-\cdots-(H_{m}\cdot C).$
Now, a proof similar to Lemma 2.8 for the surjection 3.1 yields the inequality
$\deg N_{C/X}\geq\deg Q+\deg\beta(M_{H_{1}}^{\oplus D-3})\geq\deg Q-(D-3)(H_{1}\cdot C).$
Thus,
$\deg N_{C/X}=2g(C)-2-(K_{X}\cdot C)\geq\deg Q-(D-3)(H_{1}\cdot C),$
hence,
$2g(C)-2\geq$ $(K_{X}\cdot C)-(D-3)(H_{1}\cdot C)+\deg Q$
$=$ $\sum_{i=1}^{m}(a_{i}+d_{i})(H_{i}\cdot C)-(D-3)(H_{1}\cdot C)+\deg Q.$
and by the inequality 3.2, we have
$2g(C)-2\geq\left(a_{1}+d_{1}-D+2+\frac{1}{2}\right)(H_{1}\cdot C)+\sum_{i=2}^{m}(a_{i}+d_{i}-1)(H_{i}\cdot C),$
and since, by hypothesis, $d_{i}\geq D-a_{i}-2$ for all $1\leq i\leq m$ and $D\geq 4$, we get
$2g(C)-2\geq\frac{1}{2}(H_{1}\cdot C)+\sum_{i=2}^{m}(H_{i}\cdot C).$
Therefore, in this case we have $2g(C)-2\geq\epsilon\sum_{i=1}^{m}(H_{i}\cdot C)$ for $\epsilon=\frac{1}{2}$, hence $X$ is algebraically hyperbolic.
Suppose now instead that $\Sigma\not\subset Z$. We now intersect $\Sigma$ and $Z$ to get
$([\Sigma]\cdot[Z])\cdot H_{1}=d_{1}(H_{1}\cdot C+\deg Q)+d_{2}(H_{2}\cdot C)+\cdots+d_{m}(H_{m}\cdot C).$
As $C\subset\Sigma\cap Z$, we have $([\Sigma]\cdot[Z])\cdot H_{1}\geq H_{1}\cdot C$, thus
$d_{1}(H_{1}\cdot C+\deg Q)+d_{2}(H_{2}\cdot C)+\cdots+d_{m}(H_{m}\cdot C)\geq H_{1}\cdot C,$
so
(3.3) $\deg Q\geq\left(\frac{1}{d_{1}}-1\right)(H_{1}\cdot C)-\frac{d_{2}}{d_{1}}(H_{2}\cdot C)-\cdots-\frac{d_{m}}{d_{1}}(H_{m}\cdot C).$
And as before, we have
$2g(C)-2\geq\sum_{i=1}^{m}(a_{i}+d_{i})(H_{i}\cdot C)-(D-3)(H_{1}\cdot C)+\deg Q,$
so by the inequality 3.3, we get
$2g(C)-2\geq\left(a_{1}+d_{1}-D+2+\frac{1}{d_{1}}\right)(H_{1}\cdot C)+\sum_{i=2}^{m}\left(a_{i}+d_{i}-\frac{d_{i}}{d_{1}}\right)(H_{i}\cdot C).$
Therefore, it suffices to have $a_{1}+d_{1}\geq D-2$ and $a_{i}+d_{i}-\frac{d_{i}}{d_{1}}>\epsilon$ for some $\epsilon>0$ and $2\leq i\leq m$. The first inequality is true by hypothesis. We now check the second inequality.
First, we observe that $a_{i}\leq-2$ for all $1\leq i\leq m$. Indeed, if $a_{i}\geq-1$, by Lemma 3.3, the dimension of the space of $\mathbb{P}^{N_{i}}$-lines in $A$ would be $\dim A-a_{i}-3\leq\dim A-2$, but then the lines could not cover $A$, which is not the case for a homogeneous variety (cf. [Man20, Theorem 1.5 and the following paragraph]). And since $d_{i}\geq D-a_{i}-2$, this implies that $d_{i}\geq 4$ for all $1\leq i\leq m$.
On the other hand, an $n$-dimensional projective manifold $Y$ with an ample divisor $H$ satisfies $K_{Y}\geq-(n+1)H$, with equality when $X$ is isomorphic to $\mathbb{P}^{n}$ (cf. [KO73].) Thus, looking at each projection $\pi_{i}:A\to\mathbb{P}^{N_{i}}$, we obtain that $a_{i}\geq-(D+1)$ for each $1\leq i\leq m$.
Now, rearrange the inequality we want to prove, $a_{i}+d_{i}-\frac{d_{i}}{d_{1}}>\epsilon$, as
$d_{i}>(\epsilon-a_{i})\frac{d_{1}}{d_{1}-1}.$
Since $d_{1}\geq 4$, we have $\frac{d_{1}}{d_{1}-1}\leq\frac{4}{3}$, and as $d_{i}\geq D-a_{i}-2$, it suffices to show that
$D-2-a_{i}>(\epsilon-a_{i})\frac{4}{3},$
or equivalently,
$D-2\geq-\frac{a_{i}}{3}+\epsilon^{\prime},$
for $\epsilon^{\prime}=4\epsilon/3$.
Now, as $a_{i}\geq-(D+1)$, we have $-\frac{a_{i}}{3}+\epsilon^{\prime}\leq\frac{D+1}{3}+\epsilon^{\prime}$, so we only need to show that
$D-2\geq\frac{D+1}{3}+\epsilon^{\prime}.$
Rearranging terms, this is equivalent to
$D\geq\frac{7}{2}+\frac{3\epsilon^{\prime}}{2},$
which is true since $D\geq 4$.
Therefore, in all cases we get $2g(C)-2\geq\epsilon\sum_{i=1}^{m}(H_{i}\cdot C)$ for some $\epsilon>0$, hence $X$ is algebraically hyperbolic. ∎
4. Lines in hypersurfaces of homogeneous varieties
In this section, we obtain the degrees for which the general hypersurface in $A$ contains lines. In particular, we get a bound for when $X$ is not algebraically hyperbolic. First, we collect a lemma characterizing certain Schubert classes in the Grassmannian of lines, based on Liu’s work [Liu22].
Lemma 4.1.
(cf. [Liu22, Proposition 3.1]) Let $d\geq 2$, $m>0$, and let $X\subset G(2,N)$ be a (possibly reducible) variety with class $[X]=m\cdot\sigma_{N-2,N-2-(d+1)}$ in the Chow ring of the Grassmannian of lines $G(2,N)$. Then, the lines parametrized by $X$ pass through a finite set of fixed points in $\mathbb{P}^{N-1}$.
Proof.
Consider the dual Grassmannian $G(N-2,N)$. The class of the dual variety to $[X]=m\cdot\sigma_{N-2,N-2-(d+1)}$ in $G(N-2,N)$ is $[X^{*}]=m\cdot\sigma_{2^{N-2-(d+1)},1^{d+1}}$.
By Pieri’s formula [EH16, Proposition 4.9], $[X^{*}]\cdot\sigma_{2}=0$. Since $\sigma_{2}$ parametrizes $\mathbb{P}^{N-3}$s through a point in $\mathbb{P}^{N-1}$, it follows that the general point of $\mathbb{P}^{N-1}$ is not contained in any of the $\mathbb{P}^{N-3}$s parametrized by $X^{*}$. Hence, the $\mathbb{P}^{N-3}$s parametrized by each component of $X^{*}$ sweep out a subvariety $T$ of dimension $N-2$ of $\mathbb{P}^{N-1}$. We claim that $T$ is a hyperplane, $T\cong\mathbb{P}^{N-2}$. To see this, consider the universal family
$\mathcal{U}=\{(x,p)\ |\ x\in X^{*},\ p\text{ lies in the }\mathbb{P}^{N-3}\text{ parametrized by }x\}\subset X^{*}\times T.$
The fibers of the projection $\mathcal{U}\to X^{*}$ are $\mathbb{P}^{N-3}$, so $\dim\mathcal{U}=N-3+\dim X^{*}$. Thus, by the projection $\mathcal{U}\to T$, there is a $(\dim X^{*}-1)$-parameter family of $\mathbb{P}^{N-3}$s through each point $p$ of $T$. Since $d\geq 2$, we have $\dim X^{*}-1\geq 2$, so we get a two-parameter family of $\mathbb{P}^{N-3}$s through each point $p$ of $T$. Each of these $\mathbb{P}^{N-3}$s is in the tangent space $T_{p}(T)\cong\mathbb{P}^{N-2}$, thus by dimension, $T_{p}(T)=\mathbb{P}^{N-2}=T$.
Therefore, there exists one $\mathbb{P}^{N-2}\subset\mathbb{P}^{N-1}$ such that every $\mathbb{P}^{N-3}$ parametrized by a component of $X^{*}$ is contained in $\mathbb{P}^{N-2}$. Taking back duals, every line parametrized by $X$ passes through a fixed point in $\mathbb{P}^{N-1}$. Taking the union for all components of $X$, there is a finite set of fixed points in $\mathbb{P}^{N-1}$ such that every line parametrized by $X$ passes through one of these points. ∎
Theorem 4.2.
Let $A$ be a homogeneous variety as in the Main Setting 1.3. Let $X$ be a general hypersurface of degree $(d_{1},\ldots,d_{m})$ in $A$.
  • If $d_{i}\leq\dim A-a_{i}-4$ for some $1\leq i\leq m$, then $X$ contains $\mathbb{P}^{N_{i}}$-lines. In particular, $X$ is not algebraically hyperbolic.
  • If $d_{i}>\dim A-a_{i}-4$ for some $1\leq i\leq m$, then $X$ does not contains $\mathbb{P}^{N_{i}}$-lines.
Proof.
Set $D=\dim A$. To simplify the notation, let us assume that $d_{1}\leq D-a_{1}-4$. Then, we want to show that $X$ contains $\mathbb{P}^{N_{1}}$-lines.
Since $A$ is projectively normal, there is a degree $(d_{1},\ldots,d_{m})$ hypersurface $Z\subset\mathbb{P}^{N_{1}}\times\cdots\times\mathbb{P}^{N_{m}}$ such that $X=A\cap Z$. Let $F_{1}^{N_{1}}(Z)\subset G(2,N_{1}+1)\times\mathbb{P}^{N_{2}}\times\cdots\times\mathbb{P}^{N_{m}}$ be the Fano scheme of $\mathbb{P}^{N_{1}}$-lines in $Z$, and let $F_{1}^{N_{1}}(A)$ be the Fano scheme of $\mathbb{P}^{N_{1}}$-lines in $A$. Denote by $[F_{1}^{N_{1}}(Z)]$ and $[F_{1}^{N_{1}}(A)]$ their classes in the Chow ring of $G(2,N_{1}+1)\times\mathbb{P}^{N_{2}}\times\cdots\times\mathbb{P}^{N_{m}}$. To show that $X$ contains lines, we are going to prove that the intersection class $[F_{1}^{N_{1}}(Z)]\cdot[F_{1}^{N_{1}}(A)]$ is not zero.
By the degree definition, a general $\mathbb{P}^{N_{1}}$ section of $Z$ is a general degree $d_{1}$ hypersurface in $\mathbb{P}^{N_{1}}$. By [EH16, Theorem 6.34], the scheme of lines in a general degree $d_{1}$ hypersurface has codimension $d_{1}+1$ in $G(2,N_{1}+1)$. Thus, $F_{1}^{N_{1}}(Z)$ has codimension $d_{1}+1$ in $G(2,N_{1}+1)\times\mathbb{P}^{N_{2}}\times\cdots\times\mathbb{P}^{N_{m}}$. On the other hand, by Lemma 3.3, $\dim F_{1}^{N_{1}}(A)=D-a_{1}-3$. Therefore, it suffices to prove the theorem for $D-a_{1}-3=d_{1}+1$, or equivalently, $d_{1}=D-a_{1}-4$.
The intersection of $[F_{1}^{N_{1}}(Z)]$ with $H_{2}^{N_{2}}\cdots H_{m}^{N_{m}}$ is the class of lines in a general $\mathbb{P}^{N_{1}}$ section of $Z$, which is a general degree $d_{1}$ hypersurface in $\mathbb{P}^{N_{1}}$. So, by [EH16, Proposition 6.4], we can compute the class $[F_{1}^{N_{1}}(Z)]$:
$[F_{1}^{N_{1}}(Z)]=c_{d_{1}+1}(Sym^{d_{1}}S^{*}),$
where $S$ is the universal subbundle of $G(2,N_{1}+1)$.
And since $F_{1}^{N_{1}}(A)$ is a space of $\mathbb{P}^{N_{1}}$-lines, its class $[F_{1}^{N_{1}}(A)]$ is also generated by classes of $G(2,N_{1}+1)$.
By the Whitney formula and the splitting principle [EH16, Chapter 5], we can write
$1+\sigma_{1}+\sigma_{1,1}=c(S^{*})=(1+\alpha)(1+\beta),$
for $\alpha+\beta=\sigma_{1}$ and $\alpha\beta=\sigma_{1,1}$. Thus,
$c(Sym^{d_{1}}S^{*})=(1+d_{1}\alpha)[1+(d_{1}-1)\alpha+\beta]\cdots[1+\alpha+(d_{1}-1)\beta](1+d_{1}\beta).$
We are interested in the top Chern class
$c_{d_{1}+1}(Sym^{d_{1}}S^{*})=(d_{1}\alpha)[(d_{1}-1)\alpha+\beta]\cdots[\alpha+(d_{1}-1)\beta](d_{1}\beta).$
Suppose that $d_{1}$ is even (the case when $d_{1}$ is odd is similar), and rearrange the expression above as
$[F_{1}^{N_{1}}(Z)]=c_{d_{1}+1}(Sym^{d_{1}}S^{*})$
$=(d_{1}\alpha)(d_{1}\beta)[(d_{1}-1)\alpha+\beta][\alpha+(d_{1}-1)\beta]\cdots[i\alpha+(d_{1}-i)\beta][(d_{1}-i)\alpha+i\beta]\cdots\left(\frac{d_{1}}{2}\alpha+\frac{d_{1}}{2}\beta\right)$
$=d_{1}^{2}(\alpha\beta)[(d_{1}-1)(\alpha+\beta)^{2}+(d_{1}-2)^{2}(\alpha\beta)]\cdots[i(d_{1}-i)(\alpha+\beta)^{2}+(d_{1}-2i)^{2}(\alpha\beta)]\cdots\frac{d_{1}}{2}(\alpha+\beta)$
$=d_{1}^{2}\sigma_{1,1}[(d_{1}-1)\sigma_{1}^{2}+(d_{1}-2)^{2}\sigma_{1,1}]\cdots[i(d_{1}-i)\sigma_{1}^{2}+(d_{1}-2i)^{2}\sigma_{1,1}]\cdots\frac{d_{1}}{2}\sigma_{1}.$
Notice that, after expanding terms, all coefficients are non-negative and the term $\sigma_{1,1}\sigma_{1}^{d_{1}-1}$ appears. Recall that these Schubert classes multiply by the rule: $\sigma_{1}\cdot\sigma_{k,l}=\sigma_{k+1,l}+\sigma_{k,l+1}$ and $\sigma_{1,1}\cdot\sigma_{k,l}=\sigma_{k+1,l+1}$. Thus, we can see that in the expansion of $\sigma_{1}^{d_{1}-1}$ all Schubert classes $\sigma_{i,j}$ with $i+j=d_{1}-1$ appear with positive coefficients. It follows that in the expansion of $[F_{1}^{N_{1}}(Z)]$ all classes $\sigma_{i,j}$ with $i+j=d_{1}+1$ and $i,j\geq 1$ appear with positive coefficients, that is, the only degree $d_{1}+1$ class not in the expansion of $[F_{1}^{N_{1}}(Z)]$ is $\sigma_{d_{1}+1,0}$. Since the complementary dimension class of $\sigma_{d_{1}+1,0}$ in $G(2,N+1)$ is $\sigma_{N-1,N-1-(d_{1}+1)}$, it suffices to show that $[F_{1}^{N_{1}}(A)]$ is not a multiple of $\sigma_{N-1,N-1-(d_{1}+1)}$. But, by Lemma 4.1, if $[F_{1}^{N_{1}}(A)]$ is a multiple of $\sigma_{N-1,N-1-(d_{1}+1)}$, then the lines of a general $\mathbb{P}^{N_{1}}$ section of $A$ pass through a finite set of points, which cannot happen since $A$ is homogeneous. Therefore, the product $[F_{1}^{N_{1}}(Z)]\cdot[F_{1}^{N_{1}}(A)]$ is not zero, hence there are $\mathbb{P}^{N_{1}}$-lines contained in $X$.
On the other hand, if $d_{i}>D-a_{i}-4$ for some $1\leq i\leq m$, then the dimension of $F_{1}^{N_{1}}(A)$ is less than the codimension of $F_{1}^{N_{1}}(Z)$ in $G(2,N_{i}+1)$. Since $X$ is general, $F_{1}^{N_{1}}(Z)$ is general and therefore the intersection $F_{1}^{N_{1}}(A)\cap F_{1}^{N_{1}}(Z)$ is empty. ∎
By Theorem 3.5 and Theorem 4.2, we have our main result.
Theorem 4.3.
Let $A$ be a homogeneous variety as in the Main Setting 1.3.
  • If $\dim A\geq 4$ and $d_{i}\geq\dim A-a_{i}-2$ for all $1\leq i\leq m$, then a very general hypersurface $X$ of degree $(d_{1},\ldots d_{m})$ is algebraically hyperbolic.
  • If $d_{i}\leq\dim A-a_{i}-4$ for some $1\leq i\leq m$, then a general hypersurface $X$ of degree $(d_{1},\ldots d_{m})$ contains $\mathbb{P}^{N_{i}}$-lines. In particular, $X$ is not algebraically hyperbolic.
This leaves open the cases with $d_{i}\geq\dim A-a_{i}-3$ for all $1\leq i\leq m$ and $d_{i}=\dim A-a_{i}-3$ for some $1\leq i\leq m$. When $\dim A\leq 4$, there are examples when this bound fails to imply algebraic hyperbolicity:
Example 4.4.
(cf. [Yeo25, Lemma 4.2]) A very general hypersurface $X\subset\mathbb{P}^{2}\times\mathbb{P}^{2}$ of degree $(d_{1},d_{2})$ with $d_{1}=4$ or $d_{2}=4$ contains an elliptic curve. In particular, it is not algebraically hyperbolic.
The same proof also works for $\mathbb{P}^{2}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$:
Example 4.5.
A very general hypersurface $X\subset\mathbb{P}^{2}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$ of degree $(d_{1},d_{2},d_{3})$ with $d_{1}=4$ contains an elliptic curve. In particular, it is not algebraically hyperbolic.
The bound $d_{i}\geq\dim A-a_{i}-3$ for all $i$ also fails to imply algebraic hyperbolicity in the cases $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$ and $\mathbb{P}^{2}\times\mathbb{P}^{1}$ as proved in [CR23]. For $A=\mathbb{P}^{4}$, the case of sextic threefolds remains open.
However, when $\dim A\geq 5$, it does imply algebraic hyperbolicity in the cases $A=\mathbb{P}^{n}$ and $A=\mathbb{P}^{m}\times\mathbb{P}^{n}$ as proved in [Yeo25]. We conjecture the same holds for homogeneous varieties.
Conjecture 4.6.
If $\dim A\geq 5$ and $d_{i}\geq\dim A-a_{i}-3$ for all $i$, then is $X$ algebraically hyperbolic.
5. Examples
In this section, we apply Theorem 4.3 to explicit examples of homogeneous varieties. For more details on these varieties, we refer to [Cos18]. The projective normality hypothesis follows from [RR85]. The section-dominating collections of line bundles follow from Examples 2.5 and 2.6. We remark that we can also apply the theorem to products of these varieties.
5.1. Grassmannians
Let $A=G(k,n)$ be the Grassmannian of $k$-dimensional subspaces in an $n$-dimensional vector space. Consider $G(k,n)$ embedded in $\mathbb{P}^{N}$ via the Plücker embedding $p:G(k,n)\to\mathbb{P}^{N}$. Let $H$ be the hyperplane section in $G(k,n)$ corresponding to $\mathcal{O}_{G(k,n)}(1)=p^{*}\mathcal{O}_{\mathbb{P}^{N}}(1)$. Let $\mathcal{E}$ be the line bundle $dH$ for some degree $d>0$. The dimension of the Grassmannian is $\dim G(k,n)=k(n-k)$. The tangent bundle is $T_{G(k,n)}\cong S^{*}\otimes Q$, where $S$ is the tautological bundle and $Q$ is the quotient bundle of $G(k,n)$. Then a Chern class computation gives
$c_{1}(T_{G(k,n)})=c_{1}(S^{*}\otimes Q)=nH.$
Thus, the canonical divisor is $K_{G(k,n)}=-nH$.
Theorem 5.1.
  • If $\dim G(k,n)=k(n-k)\geq 4$ and $d\geq k(n-k)+n-2$, then a very general degree $d$ hypersurface $X\subset G(k,n)$ is algebraically hyperbolic.
  • If $d\leq k(n-k)+n-4$, then a general degree $d$ hypersurface $X$ contains a line. In particular, $X$ is not algebraically hyperbolic.
5.2. Products of Grassmannians
Let $A$ be the product of Grassmannians $A=\prod_{i=1}^{m}G(k_{i},n_{i})$. We can see $A\subset\mathbb{P}^{N_{1}}\times\cdots\times\mathbb{P}^{N_{m}}$ via the Plücker embedding for each factor. Let $H_{i}$ be the pullback of the hyperplane section via each projection $A\to G(k_{i},n_{i})$. For $d_{i}\geq 0$, let $\mathcal{E}=\sum_{i=1}^{m}d_{i}H_{i}$. The dimension of $A$ is $\sum_{i=1}^{m}k_{i}(n_{i}-k_{i})$. The canonical divisor of $A$ is $K_{A}=\sum_{i=1}^{m}(-n_{i})H_{i}$.
Theorem 5.2.
(Products of Grassmannians)
  • If $\dim A=\sum_{i=1}^{m}k_{i}(n_{i}-k_{i})\geq 4$ and $d_{i}\geq\left(\sum_{j=1}^{m}k_{j}(n_{j}-k_{j})\right)+n_{i}-2$, for all $1\leq i\leq m$, then a very general hypersurface of $\prod_{i=1}^{m}G(k_{i},n_{i})$ of degree $(d_{1},\ldots,d_{m})$ is algebraically hyperbolic.
  • If $d_{i}\leq\left(\sum_{j=1}^{m}k_{j}(n_{j}-k_{j})\right)+n_{i}-4$ for some $1\leq i\leq m$, then a general hypersurface of degree $(d_{1},\ldots,d_{m})$ of $\prod_{i=1}^{m}G(k_{i},n_{i})$ contains a line. In particular, it is not algebraically hyperbolic.
5.3. Orthogonal Grassmannians
Let $A$ be the orthogonal Grassmannian $OG(k,n)$. Similarly, let $H$ be the hyperplane section, and $\mathcal{E}=dH$ for $d>0$. We have $\dim OG(k,n)=\frac{k(2n-3k-1)}{2}$. The orthogonal Grassmannian is the zero locus in $G(k,n)$ of a global section of $\operatorname{Sym}^{2}(S^{*})$, and we have an exact sequence of tangent bundles
$0\longrightarrow T_{OG(k,n)}\longrightarrow S^{*}\otimes Q\longrightarrow\operatorname{Sym}^{2}(S^{*})\longrightarrow 0$
So, by a Chern class computation, we have
$c_{1}(T_{OG(k,n)})=c_{1}(S^{*}\otimes Q)-c_{1}(\operatorname{Sym}^{2}(S^{*}))=(n-k-1)H.$
Thus, the canonical divisor of $OG(k,n)$ is $K_{OG}=-(n-k-1)H$.
Theorem 5.3.
(Orthogonal Grassmannians)
  • If $\dim OG(k,n)=\frac{k(2n-3k-1)}{2}\geq 4$ and $d\geq\frac{k(2n-3k-1)}{2}+n-k-3$, then a very general hypersurface of $OG(k,n)$ of degree $d$ is algebraically hyperbolic.
  • If $d\leq\frac{k(2n-3k-1)}{2}+n-k-5$, then a general hypersurface of degree $d$ of $OG(k,n)$ contains a line. In particular, it is not algebraically hyperbolic.
5.4. Symplectic Grassmannians
Let $A=SG(k,n)$ be a symplectic Grassmannian. Let $H$ by the hyperplane section, and $\mathcal{E}=dH$ for $d>0$. In this case, $\dim SG(k,n)=\frac{k(2n-3k+1)}{2}$. Similarly, we have an exact sequence
$0\longrightarrow T_{SG(k,n)}\longrightarrow S^{*}\otimes Q\longrightarrow\wedge^{2}(S^{*})\longrightarrow 0,$
so we can compute
$c_{1}(T_{SG(k,n)})=c_{1}(S^{*}\otimes Q)-c_{1}(\wedge^{2}(S^{*}))=(n-k+1)H.$
Thus, the canonical divisor is $K_{SG(k,n)}=-(n-k+1)H$.
Theorem 5.4.
(Symplectic Grassmannians)
  • If $\dim A=\frac{k(2n-3k+1)}{2}\geq 4$ and $d\geq\frac{k(2n-3k+1)}{2}+n-k-1$, then a very general degree $d$ hypersurface of $SG(k,n)$ is algebraically hyperbolic.
  • If $d\leq\frac{k(2n-3k+1)}{2}+n-k-3$, then a general degree $d$ hypersurface of $SG(k,n)$ contains a line. In particular, it is not algebraically hyperbolic.
5.5. Flag varieties
Let $A=F(k_{1},\ldots,k_{m};n)$ be a flag variety. The projections
$F(k_{1},\ldots,k_{m};n)\to G(k_{1},n)\times\cdots\times G(k_{m},n)$
give, via pull-back, the hyperplane classes $H_{1},\ldots,H_{m}$ that generate the Picard group of $A$. To fix notation, let $k_{0}=0$ and $k_{m+1}=n$. The dimension of $A$ is $\sum_{i=1}^{m}k_{i}(k_{i+1}-k_{i})$. The tangent bundle of $A$ is isomorphic to $\bigoplus_{i=1}^{m}S_{k_{i}}^{*}\otimes(S_{k_{i+1}}/S_{k_{i}})$, where $S_{k_{i}}$ is the rank $k_{i}$ tautological bundle. Then, by a Chern class computation,
$c_{1}(F(k_{1},\ldots,k_{m};n))=c_{1}\left(\bigoplus_{i=1}^{m}S_{k_{i}}^{*}\otimes(S_{k_{i+1}}/S_{k_{i}})\right)=\sum_{i=1}^{m}(k_{i+1}-k_{i-1})H_{i}.$
Therefore, the canonical divisor of $F(k_{1},\ldots,k_{m};n)$ is $K_{A}=-\sum_{i=1}^{m}(k_{i+1}-k_{i-1})H_{i}=-k_{2}H_{1}-(k_{3}-k_{1})H_{2}-\cdots-(n-k_{m-1})H_{m}$.
Theorem 5.5.
(Flag varieties)
  • If $\dim A=\sum_{i=1}^{m}k_{i}(k_{i+1}-k_{i})\geq 4$ and $d_{i}\geq\left(\sum_{j=1}^{m}k_{j}(k_{j+1}-k_{j})\right)+k_{i+1}-k_{i-1}-2$ for all $1\leq i\leq m$, then a very general hypersurface of $F(k_{1},\ldots,k_{m};n)$ of degree $(d_{1},\ldots,d_{m})$ is algebraically hyperbolic.
  • If $d_{i}\leq\left(\sum_{j=1}^{m}k_{j}(k_{j+1}-k_{j})\right)+k_{i+1}-k_{i-1}-4$ for some $i$, then a general hypersurface of $F(k_{1},\ldots,k_{m};n)$ of degree $(d_{1},\ldots,d_{m})$ contains a line. In particular, it is not algebraically hyperbolic.